# Nonreciprocal Transmission in Hybrid Atomic Ensemble-Optomechanical Systems

### v1

@@ Abstract
We investigate perfect optical nonreciprocal transmission in a hybrid optomechanical system that incorporates an atomic ensemble. By introducing complex coupling strengths between the atomic ensemble and a mechanical oscillator, nonreciprocity is induced through interference between distinct optical pathways. The nonreciprocal transmission is governed by the real and imaginary components of the coupling constants, along with the relative phase differences between the optomechanical couplings. Our analysis reveals that, with precise tuning of system parameters, such as coupling strengths, detuning, and phase differences, perfect nonreciprocity can be achieved. We derive the conditions necessary for optimal nonreciprocal transmission and demonstrate its dependence on the complex nature of the coupling. These findings offer valuable insights for designing nonreciprocal optical devices, including isolators and circulators, with potential applications in quantum communication, signal processing, and photonics.
@@ I Introduction
Interaction between atomic systems and mechanical oscillators has garnered considerable attention due to its potential in probing quantum phenomena and enabling novel applications. When coupled through radiation pressure, the mechanical system interacts with an optical field, constituting a cavity optomechanical (COM) system. These systems provide a versatile platform for exploring light-matter interactions at the quantum level. Notably, trapped atomic ensembles in optical cavities, which are strongly coupled to the optical field and mechanical oscillators, have demonstrated significant effects, such as quantum correlations and entanglement [1, 2, 3, 4, 5, 6]. These hybrid systems hold considerable promise for applications in quantum information processing [7, 8], mass sensing [9, 10, 11], fundamental investigations of quantum mechanics at macroscopic scales [12, 13], and nonlinear dynamics [14, 15].
@@
A key functionality for quantum technologies is nonreciprocity, which facilitates unidirectional signal transmission and is pivotal in devices such as optical isolators and circulators. Such nonreciprocal devices are required for preventing undesired reflections and safeguarding sensitive optical components, such as lasers, from backscatter-induced instabilities [16, 17]. The realization of nonreciprocal transmission requires the breaking of time-reversal symmetry, a condition that has been theoretically and experimentally achieved through various mechanisms, including magnetic fields, optomechanical interactions, and engineered dissipative processes [18, 19, 20, 21, 22].
@@
Recent advances in cavity optomechanics have unveiled new pathways for achieving nonreciprocal optical behavior without the reliance on magnetic effects. Specifically, robust optomechanical coupling has been demonstrated to induce nonreciprocal transparency and phase shifts in microring resonators [23]. Additionally, non-Hermitian physics has emerged as a potent framework for describing open quantum systems, wherein energy dissipation or gain plays a fundamental role. Non-Hermitian Hamiltonians, which incorporate complex coupling terms, have been extensively used to model nonreciprocal transmission in optical and mechanical systems, leveraging the interplay between loss and interference [24, 25, 20].
@@
In this work, we investigate optical nonreciprocity in a hybrid atomic ensemble-cavity optomechanical system. By adjusting the system parameters—particularly the complex coupling strength $J_{3}$, which models dissipative interactions between the atomic ensemble and the mechanical mode—perfect nonreciprocity can be achieved. Our methodology employs a non-Hermitian Hamiltonian to describe the system, with its imaginary component $J_{3}$ symbolizing energy dissipation, resulting in unidirectional signal transmission. This is consistent with recent advances in non-Hermitian physics, where dissipation-induced nonreciprocity has been experimentally demonstrated in optical cavities and waveguides [20, 21].
@@
The structure of this paper is as follows: In Section II, we introduce the theoretical model of the hybrid atomic ensemble-cavity optomechanical system and derive its Hamiltonian and equations of motion. In Section III, we analyze the conditions required for nonreciprocal transmission and study the effects of system parameters on the transmission properties. Finally, we conclude with a summary of our findings and discuss potential applications in quantum information processing and photonics.
@@ II.1 Model
We consider a hybrid optomechanical system consisting of two coupled cavities, one of which contains a trapped atomic ensemble, as illustrated in Figure 1. The first cavity mode is coupled to the atomic ensemble, while the two cavities are connected via photon tunneling. A mechanical oscillator is coupled to one of the cavities through radiation pressure, enabling interactions between the optical field and mechanical motion. The system is driven by strong external fields and probed by weak fields to analyze its optical response.
Figure 1: (a) A hybrid optomechanical system with atomic ensemble in one cavity. The coupling fields (probe fields) with amplitudes $\mathcal{E}_{1}$ and $\mathcal{E}_{2}~{}(\mathcal{E}_{p1}$ and $\mathcal{E}_{p2})$ are used to drive the two cavities that are coupled to each other via radiation pressure. (b) The diagram of interactions among subsystems in the optomechanical system.
@@
where $a_{j}(a_{j}^{\dagger})$ and $b(b^{\dagger})$ are the annihilation (creation) operators for the cavity and mechanical modes (with $j=1,2$), obeying the commutation relations $[a_{j},a_{j}^{\dagger}]=1$ and $[b,b^{\dagger}]=1$. The atomic ensemble, modeled as a bosonic mode, interacts with both the optical and mechanical subsystems. Its Hamiltonian is written as:
$H_{en}$ $=\frac{\Delta_{en}}{2}\sum_{j=1}^{N}\sigma_{j}^{z}+\left(J_{2}^{\ast}a_{1}\sum_{j=1}^{N}\sigma_{j}^{+}+J_{2}a_{1}^{\dagger}\sum_{j=1}^{N}\sigma_{j}^{-}\right)$ (3)
$+J_{3}\left(\sum_{j=1}^{N}\sigma_{j}^{+}+\sum_{j=1}^{N}\sigma_{j}^{-}\right)(b+b^{\dagger}),$
where $J_{2}$ is the coupling strength between the atomic ensemble and the first cavity mode, and $J_{3}$ describes the coupling between the atomic ensemble and the mechanical oscillator. The detuning of the atomic ensemble mode from the driving frequency is defined as $\Delta_{en}=\omega_{en}-\omega_{L}$. We assume that the number of atoms in the excited state is significantly lower than the total number of atoms $N$, indicating that the atomic ensemble operates under low excitation conditions. Consequently, we can treat the atomic ensemble as a bosonic mode. The atomic ensemble is initially modeled using spin-1/2 Pauli operators $\sigma_{j}^{z}$ and $\sigma_{j}^{\pm}$ for the $j^{th}$ two-level atom, which satisfy the commutation relations $[\sigma_{j}^{+},\sigma_{j}^{-}]=\sigma_{j}^{z}$ and $[\sigma_{j}^{z},\sigma_{j}^{\pm}]=\pm 2\sigma_{j}^{\pm}$. To treat the atomic ensemble as a bosonic mode, we apply the Holstein-Primakoff transformation, which enables us to map the collective spin operators to bosonic creation and annihilation operators,
@@
$H_{en}=\Delta_{en}d^{\dagger}d+\left(J_{2}^{\ast}a_{1}d^{\dagger}+J_{2}a_{1}^{\dagger}d\right)+J_{3}(d+d^{\dagger})(b+b^{\dagger}).$ (4)
@@
$H_{int}=g_{1}a_{1}^{\dagger}a_{1}(b+b^{\dagger})+g_{2}a_{2}^{\dagger}a_{2}(b+b^{\dagger})+J_{1}(a_{1}a_{2}^{\dagger}+a_{2}a_{1}^{\dagger}),$ (5)
@@
$\begin{split}H_{driv}&=i(\mathcal{E}_{1}a_{1}^{\dagger}-\mathcal{E}_{1}^{\ast}a_{1})+i(\mathcal{E}_{2}a_{2}^{\dagger}-\mathcal{E}_{2}^{\ast}a_{2})\\ &+i\mathcal{E}_{p1}(a_{1}^{\dagger}e^{-i\delta t}-a_{1}e^{i\delta t})+i\mathcal{E}_{p2}(a_{2}^{\dagger}e^{-i\delta t}-a_{2}e^{i\delta t}),\end{split}$ (6)
@@
The complex coupling strength $J_{3}$ plays a critical role in inducing nonreciprocal transmission. The real part of $J_{3}$ governs the interaction strength between the atomic ensemble and the mechanical oscillator, while the imaginary component represents dissipative processes, where energy is either lost to the environment or exchanged between subsystems. The introduction of dissipation through the imaginary part of $J_{3}$ breaks the symmetry of the system, leading to unidirectional signal transmission. This dissipation-induced nonreciprocity is a hallmark of non-Hermitian systems, which allows for nonreciprocal effects without the need for magnetic fields. For experimental realization of this non-Hermitian Hamiltonian, careful consideration must be given to coupling between atomic ensembles and mechanical modes. This can be achieved through engineered dissipative processes by tuning interaction strengths and decay rates in optomechanical systems with external baths.
@@ II.2 Dynamical Equations
$\begin{cases}(i\Delta^{\prime}_{1}+\kappa_{1})\alpha_{1}+iJ_{1}\alpha_{2}+iJ_{2}\rho-\mathcal{E}_{1}=0,\\ (i\Delta^{\prime}_{2}+\kappa_{2})\alpha_{2}+iJ_{1}\alpha_{1}-\mathcal{E}_{2}=0,\\ (i\Delta_{en}+f)\rho+iJ^{\ast}_{2}\alpha_{1}+2iJ_{3}\Re(\beta)=0,\\ (i\omega_{m}+\gamma)\beta+ig_{1}|\alpha_{1}|^{2}+ig_{2}|\alpha_{2}|^{2}+2iJ_{3}\Re(\rho)=0,\end{cases}$ (9)
@@
$\begin{cases}\delta\dot{a_{1}}=-(i\Delta^{\prime}_{1}+\kappa_{1})\delta a_{1}-i|J_{2}|e^{i\phi}\delta d-iJ_{1}\delta a_{2}\\ -iG_{1}(\delta b+\delta b^{\dagger})+\mathcal{E}_{p1}e^{-i\delta t},\\ \delta\dot{a_{2}}=-(i\Delta^{\prime}_{2}+\kappa_{2})\delta a_{2}-iJ_{1}\delta a_{1}-iG_{2}e^{i\theta}(\delta b+\delta b^{\dagger})\\ +\mathcal{E}_{p2}e^{-i\delta t},\\ \delta\dot{d}=-(i\Delta_{en}+f)\delta d-i|J_{2}|e^{-i\phi}\delta a_{1}-iJ_{3}(\delta b+\delta b^{\dagger}),\\ \delta\dot{b}=-(i\omega_{m}+\gamma)\delta b-iG_{1}(\delta a_{1}+\delta a^{\dagger}_{1})-iG_{2}(e^{-i\theta}\delta a_{2}\\ +e^{i\theta}\delta a^{\dagger}_{2})-iJ_{3}(\delta d+\delta d^{\dagger}),\end{cases}$ (10)
where the effective optomechanical coupling strengths are defined as $G_{1}=g_{1}\alpha_{1}$ and $G_{2}=g_{2}\alpha_{2}e^{-i\theta}$. The phase difference between the effective optomechanical couplings is $\theta=\theta_{2}-\theta_{1}$. Without loss of generality, we assume that $G_{1}$ is real, i.e., $\theta_{1}=0$, which reduces the phase difference to $\theta=\theta_{2}$. The quantum Langevin equations in Eq.(10) describe the small fluctuations around the steady-state values and govern the quantum dynamics of the system. These linearized equations incorporate the coupling between the optical, mechanical, and atomic subsystems, as well as dissipation via the terms involving $\kappa_{j}$, $\gamma$, and $f$. The coupling terms involving $J_{2}$ and $J_{3}$ play a central role in generating nonreciprocity, with the dissipation modeled by the imaginary part of $J_{3}$ leading to asymmetric transmission. The full quantum dynamics, including these fluctuations, can be used to calculate observables such as the transmission coefficients, which reveal the nonreciprocal behavior of the system. These calculations are essential for understanding how energy dissipation and interference in the system give rise to unidirectional optical transmission. To solve Eq. (10), we perform the transformations $\delta a_{j}\rightarrow\delta\tilde{a}_{j}e^{-i\Delta^{\prime}_{j}t}$, $\delta d\rightarrow\delta\tilde{d}e^{-i\Delta_{en}t}$, and $\delta b\rightarrow\delta\tilde{b}e^{-i\omega_{m}t}$. These transformations yield,

### v2

@@ Abstract
We investigate perfect optical nonreciprocal transmission in a hybrid optomechanical system that incorporates an atomic ensemble. By introducing complex coupling strengths between the atomic ensemble and a mechanical oscillator, nonreciprocity is induced through interference between distinct optical pathways. The nonreciprocal transmission is governed by the real and imaginary components of the coupling constants, along with the relative phase differences between the optomechanical couplings. Our analysis reveals that, with precise tuning of system parameters, such as coupling strengths, detuning, and phase differences, perfect nonreciprocity can be achieved. We derive the conditions necessary for optimal nonreciprocal transmission and demonstrate its dependence on the complex nature of the coupling. These findings offer valuable insights for the design of nonreciprocal optical devices, including isolators and circulators, with potential applications in quantum communication, signal processing, and photonics.
@@ I Introduction
The interaction between atomic systems and mechanical oscillators has garnered considerable attention because of its potential in probing quantum phenomena and enabling novel applications. When coupled through radiation pressure, the mechanical system interacts with an optical field, which constitutes a cavity optomechanical (COM) system. These systems provide a versatile platform for exploring light-matter interactions at the quantum level. In particular, atomic ensembles trapped in optical cavities, which are strongly coupled to the optical field and mechanical oscillators, have demonstrated significant effects, such as quantum correlations and entanglement [1, 2, 3, 4, 5, 6]. These hybrid systems hold considerable promise for applications in quantum information processing [7, 8], mass sensing [9, 10, 11], fundamental investigations of quantum mechanics at macroscopic scales [12, 13], and nonlinear dynamics [14, 15].
@@
A key functionality for quantum technologies is nonreciprocity, which facilitates unidirectional signal transmission and is pivotal in devices such as optical isolators and circulators. Such nonreciprocal devices are required to prevent undesired reflections and protect sensitive optical components, such as lasers, from backscatter-induced instabilities [16, 17]. The realization of nonreciprocal transmission requires breaking of time-reversal symmetry, a condition that has been achieved theoretically and experimentally through various mechanisms, including magnetic fields, optomechanical interactions, and engineered dissipative processes [18, 19, 20, 21, 22].
@@
Recent advances in cavity optomechanics have unveiled new pathways for achieving nonreciprocal optical behavior without the reliance on magnetic effects. Specifically, robust optomechanical coupling has been shown to induce nonreciprocal transparency and phase shifts in microring resonators [23]. Additionally, the Lindblad master equation provides a rigorous framework for describing the dynamics of open quantum systems. Account for both unitary evolution and the effects of the external environment. Non-Hermitian Hamiltonians, which incorporate complex coupling terms, have been widely used to model nonreciprocal transmission in optical and mechanical systems, utilizing the interaction between loss and interference [24, 25, 20].
@@
In this work, we investigate optical nonreciprocity in a hybrid atomic ensemble-cavity optomechanical system. By adjusting the system parameters—particularly the complex coupling strength $J_{3}$, which models dissipative interactions between the atomic ensemble and the mechanical mode—perfect nonreciprocity can be achieved. Our methodology employs a non-Hermitian Hamiltonian to describe the system, with its imaginary component $J_{3}$ symbolizing energy dissipation, resulting in transmission of unidirectional signals. This is consistent with recent advances in non-Hermitian physics, where dissipation-induced nonreciprocity has been experimentally demonstrated in optical cavities and waveguides [20, 21].
@@
The structure of this paper is as follows: In Section II, we introduce the theoretical model of the hybrid atomic ensemble-cavity optomechanical system and derive its Hamiltonian and equations of motion. In Section III, we analyze the conditions required for nonreciprocal transmission and study the effects of system parameters on transmission properties. Finally, we conclude with a summary of our findings and discuss potential applications in quantum information processing and photonics.
@@ II.1 Model
We consider a hybrid optomechanical system consisting of two coupled cavities, one of which contains a trapped atomic ensemble, as illustrated in Figure 1. The first cavity mode is coupled to the atomic ensemble, whereas the two cavities are connected via photon tunneling. A mechanical oscillator is coupled to both cavities through radiation pressure, enabling interactions between the optical field and mechanical motion. The system is driven by strong external fields and probed by weak fields to analyze its optical response.
Figure 1: (a) A hybrid optomechanical system with atomic ensemble in one cavity. The coupling fields (probe fields) with amplitudes $\mathcal{E}_{1}$ and $\mathcal{E}_{2}~{}(\mathcal{E}_{p1}$ and $\mathcal{E}_{p2})$ are used to drive the two cavities that are tunnel-coupled and coupled to the common mechanical mode via radiation pressure. (b) The diagram of interactions among subsystems in the optomechanical system.
@@
where $a_{j}(a_{j}^{\dagger})$ and $b(b^{\dagger})$ are the annihilation (creation) operators for the cavity and mechanical modes (with $j=1,2$), obeying the commutation relations $[a_{j},a_{j}^{\dagger}]=1$ and $[b,b^{\dagger}]=1$. We assume that the number of atoms in the excited state is significantly lower than the total number of atoms $N$, indicating that the atomic ensemble operates under low excitation conditions. Consequently, we can treat the atomic ensemble as a bosonic mode. The atomic ensemble is initially modeled using spin-1/2 Pauli operators $\sigma_{j}^{z}$ and $\sigma_{j}^{\pm}$ for the $j^{th}$ two-level atom, which satisfy the commutation relations $[\sigma_{j}^{+},\sigma_{j}^{-}]=\sigma_{j}^{z}$ and $[\sigma_{j}^{z},\sigma_{j}^{\pm}]=\pm 2\sigma_{j}^{\pm}$. To treat the atomic ensemble as a bosonic mode, we apply the Holstein-Primakoff transformation, which enables us to map the collective spin operators to bosonic creation and annihilation operators,
@@
The atomic ensemble interacts with both the optical and mechanical subsystems. Its Hamiltonian is written as:
$H_{\rm en}$ $=\frac{\Delta_{\rm en}}{2}\sum_{j=1}^{N}\sigma_{j}^{z}+\left(J_{2}^{\ast}a_{1}\sum_{j=1}^{N}\sigma_{j}^{+}+J_{2}a_{1}^{\dagger}\sum_{j=1}^{N}\sigma_{j}^{-}\right)$ (3)
$+J_{m}\left(\sum_{j=1}^{N}\sigma_{j}^{+}+\sum_{j=1}^{N}\sigma_{j}^{-}\right)(b+b^{\dagger}),$
where $J_{2}$ is the coupling strength between the atomic ensemble and the first cavity mode, and $J_{m}$ describes the coupling between the atomic ensemble and the mechanical oscillator. The detuning of the atomic ensemble mode from the driving frequency is defined as $\Delta_{\rm en}=\omega_{\rm en}-\omega_{L}$. The coupling $J_{m}$ is usually considered to be real in most quantum physics context. However, in systems that interact with an external environment such as bath, the coupling can become complex due to non-hermitian dynamics which include dissipation (loss) or amplification (gain). The system under consideration is model as an open quantum system where loss or gain is induced. The Hamiltonian of this type of system can acquire complex eigenvalues which represent the system that is not in equilibrium. The complex eigenvalues indicate that the Hamiltonian is non-hermitian, i.e., $H_{\rm com}\not=H_{\rm com}^{\dagger}$. Non-hermitian Hamiltonian usually have complex eigenvalues, where the imaginary part relates system’s dissipation or amplification. The dynamics of the density matrix $\rho$ for an open quantum system is described by the Lindblad master equation,
$\dot{\rho}=-i[H_{\rm com},\rho]+\mathcal{D}[\rho],$ (4)
where $\mathcal{D}[\rho]$ is the dissipator term that account for the interaction with the environment,
$\mathcal{D}[\rho]=\sum_{j=1}\Gamma_{j}\left(L_{j}\rho L_{j}^{\dagger}-\frac{1}{2}\{L_{j}^{\dagger}L_{j},\rho\}\right).$ (5)
In this work, our mechanical oscillator can lose energy to the environment or gain energy from the external laser field since it is interacting with the cavity. Because of the gain and loss induce in the system, $J_{m}$ can be treated as if had imaginary part to qualitatively describe these processes. In this light, $J_{m}$ takes the form $J_{m}=\text{Re}[J_{m}]+i\text{Im}[J_{m}]$. The Hamiltonian defining the interaction between atomic ensemble and mechanical oscillator takes the form
$H_{\rm com}=\left(\text{Re}[J_{m}]+i\text{Im}[J_{m}]\right)\left(\sum_{j=1}^{N}\sigma^{+}_{j}+\sum_{j=1}^{N}\sigma^{-}_{j}\right)(b+b^{\dagger}),$ (6)
from the commutator relation,
$\begin{split}-i[H_{\rm com},\rho]=&-i\Big{[}\left(\text{Re}[J_{m}]+i\text{Im}[J_{m}]\right)\Big{.}\\ &\Big{.}\left(\sum_{j=1}^{N}\sigma^{+}_{j}+\sum_{j=1}^{N}\sigma^{-}_{j}\right)(b+b^{\dagger}),\rho\Big{]}.\end{split}$ (7)
Based on this expression, we have two contributions, one is from the real part and the other one is from the imaginary part respectively, represented as:
$-i\left[\left(\text{Re}[J_{3}]\right)\left(\sum_{j=1}^{N}\sigma^{+}_{j}+\sum_{j=1}^{N}\sigma^{-}_{j}\right)(b+b^{\dagger}),\rho\right],$ (8)
$-i\left[\left(i\text{Im}[J_{3}]\right)\left(\sum_{j=1}^{N}\sigma^{+}_{j}+\sum_{j=1}^{N}\sigma^{-}_{j}\right)(b+b^{\dagger}),\rho\right].$ (9)
In the dissipative regime, that is, where the dissipative processes are dominant, we can neglect the real part of $J_{m}$ in order to emphasize on the effects of loss and gain. The last term in Eq. (3) now becomes,
$H_{com}=iJ_{m}\left(\sum_{j=1}^{N}\sigma^{+}_{j}+\sum_{j=1}^{N}\sigma^{-}_{j}\right)(b+b^{\dagger}).$ (10)
For simplicity, we define $J_{3}=iJ_{m}$, this purely imaginary coupling renders the Hermitian Hamiltonian above non-Hermitian which is indicative of open quantum system exhibiting gain or loss. Therefore, we can consider the parameter $J_{3}$ as the degree of non-Hermiticity. The imaginary part of the coupling can introduce effective decay or amplification in the system’s dynamics influencing the evolution of the density matrix through the dissipative term $\mathcal{D}[\rho]$. We note that the complex coupling $J_{3}$ is introduced as an effective parameter to capture the dissipative dynamics between the atomic ensemble and the mechanical oscillator in a non-Hermitian framework. This approach approximates the effects of environmental interactions, such as energy loss or gain, without explicitly modeling the bath dynamics through a full Lindblad master equation. In a complete Lindblad picture, the dissipative coupling would arise from the interplay between the Hermitian interaction Hamiltonian and the Lindblad dissipator terms, which describe specific dissipative channels (e.g., atomic decay or mechanical damping). The imaginary part of $J_{3}$ qualitatively represents the asymmetric energy exchange induced by the bath, which can be engineered, for example, by tuning the decay rates of atomic transitions using external laser fields. While our non-Hermitian treatment simplifies the analysis and highlights the role of dissipation in inducing nonreciprocity, it does not account for stochastic quantum jumps or the detailed bath structure, which would be captured in a full Lindblad formalism. This approximation is justified for studying the qualitative features of nonreciprocal transmission, as supported by similar approaches in the literature (e.g., Ref. [26]).
@@
$H_{\rm en}=\Delta_{\rm en}d^{\dagger}d+\left(J_{2}^{\ast}a_{1}d^{\dagger}+J_{2}a_{1}^{\dagger}d\right)+J_{3}(d+d^{\dagger})(b+b^{\dagger}).$ (11)
To provide insight into the microscopic origin of the complex coupling $J_{3}$, consider a scenario where the atomic ensemble and mechanical oscillator are coupled to a common bath, such as a thermal reservoir or an engineered environment. The Hermitian interaction between the atomic ensemble and the mechanical oscillator, typically of the form $H_{\text{\rm int}}=g(d+d^{\dagger})(b+b^{\dagger})$, is modified by the bath-induced dissipation. By adiabatically eliminating the bath degrees of freedom using the Born-Markov approximation, the effective interaction can acquire an imaginary component due to correlated dissipation or phase-dependent interference between the coherent coupling and dissipative channels. For instance, external laser fields can be used to control the atomic decay rates, introducing an effective dissipative coupling that renders the Hamiltonian non-Hermitian. This mechanism is analogous to the imaginary coupling between optical resonators induced by gain or loss, as demonstrated by Takata et al. [26]. Our phenomenological use of $J_{3}$ captures these effects in a simplified manner, enabling us to focus on the resulting nonreciprocal transmission without solving the full bath dynamics.
@@
$H_{\rm int}=g_{1}a_{1}^{\dagger}a_{1}(b+b^{\dagger})+g_{2}a_{2}^{\dagger}a_{2}(b+b^{\dagger})+J_{1}(a_{1}a_{2}^{\dagger}+a_{2}a_{1}^{\dagger}),$ (12)
@@
$\begin{split}H_{\rm driv}&=i(\mathcal{E}_{1}a_{1}^{\dagger}-\mathcal{E}_{1}^{\ast}a_{1})+i(\mathcal{E}_{2}a_{2}^{\dagger}-\mathcal{E}_{2}^{\ast}a_{2})\\ &+i\mathcal{E}_{p1}(a_{1}^{\dagger}e^{-i\delta t}-a_{1}e^{i\delta t})+i\mathcal{E}_{p2}(a_{2}^{\dagger}e^{-i\delta t}-a_{2}e^{i\delta t}),\end{split}$ (13)
@@
The complex coupling strength $J_{3}$ plays a critical role in inducing nonreciprocal transmission. The real part of $J_{3}$ governs the interaction strength between the atomic ensemble and the mechanical oscillator, while the imaginary component represents dissipative processes, where energy is either lost to the environment or exchanged between subsystems. The introduction of dissipation through the imaginary part of $J_{3}$ breaks the symmetry of the system, leading to unidirectional signal transmission. The dissipation-induced nonreciprocity is a characteristic of non-Hermitian systems which can lead to nonreciprocal effects without the for traditionally used magnetic fields, although magnetic fields still play a vital role in certain systems to achieve nonreciprocal behaviour. A microscopic prescription involves coupling an atomic cloud for example, cold atoms or Bose-Einstein condensate to a mechanical oscillator such as a membrane using external laser fields. Dissipative mechanism is introduce into a cavity where mechanical oscillator and atomic ensemble are interacting via interaction with the bath (environment). By tuning the decay rate of the atomic transitions using external fields, one can introduce an effective dissipative coupling with an imaginary term. This renders the Hamiltonian non-hermitian and induces loss or gain into the system.
In Ref. [26], the concept of imaginary coupling in non-Hermitian coupled-mode is explicitly demonstrated. The authors successfully studied the imaginary coupling between optical resonators induced by gain or loss. This study provides a theoretical framework for realizing complex couplings in quantum systems, which we extend to our optomechanical setup with an atomic ensemble. In our model, the imaginary coupling $J_{3}$ is introduced phenomenologically to capture similar dissipative effects, approximating the complex dynamics that would arise from a full Lindblad treatment of the system-bath interaction.
@@ II.2 Dynamical equations
$\begin{cases}(i\Delta^{\prime}_{1}+\kappa_{1})\alpha_{1}+iJ_{1}\alpha_{2}+iJ_{2}\rho-\mathcal{E}_{1}=0,\\ (i\Delta^{\prime}_{2}+\kappa_{2})\alpha_{2}+iJ_{1}\alpha_{1}-\mathcal{E}_{2}=0,\\ (i\Delta_{\rm en}+f)\rho+iJ^{\ast}_{2}\alpha_{1}+2iJ_{3}\Re(\beta)=0,\\ (i\omega_{m}+\gamma)\beta+ig_{1}|\alpha_{1}|^{2}+ig_{2}|\alpha_{2}|^{2}+2iJ_{3}\Re(\rho)=0,\end{cases}$ (16)
@@
$\begin{cases}\delta\dot{a_{1}}=-(i\Delta^{\prime}_{1}+\kappa_{1})\delta a_{1}-i|J_{2}|e^{i\phi}\delta d-iJ_{1}\delta a_{2}\\ -iG_{1}(\delta b+\delta b^{\dagger})+\mathcal{E}_{p1}e^{-i\delta t},\\ \delta\dot{a_{2}}=-(i\Delta^{\prime}_{2}+\kappa_{2})\delta a_{2}-iJ_{1}\delta a_{1}-iG_{2}e^{i\theta}(\delta b+\delta b^{\dagger})\\ +\mathcal{E}_{p2}e^{-i\delta t},\\ \delta\dot{d}=-(i\Delta_{\rm en}+f)\delta d-i|J_{2}|e^{-i\phi}\delta a_{1}-iJ_{3}(\delta b+\delta b^{\dagger}),\\ \delta\dot{b}=-(i\omega_{m}+\gamma)\delta b-iG_{1}(\delta a_{1}+\delta a^{\dagger}_{1})-iG_{2}(e^{-i\theta}\delta a_{2}\\ +e^{i\theta}\delta a^{\dagger}_{2})-iJ_{3}(\delta d+\delta d^{\dagger}),\end{cases}$ (17)
where the effective optomechanical coupling strengths are defined as $G_{1}=g_{1}\alpha_{1}$ and $G_{2}=g_{2}\alpha_{2}e^{-i\theta}$. The phase difference between the effective optomechanical couplings is $\theta=\theta_{2}-\theta_{1}$. Without loss of generality, we assume that $G_{1}$ is real, i.e., $\theta_{1}=0$, which reduces the phase difference to $\theta=\theta_{2}$. The quantum Langevin equations in Eq.(17) describe the small fluctuations around the steady-state values and govern the quantum dynamics of the system. These linearized equations incorporate the coupling between the optical, mechanical, and atomic subsystems, as well as dissipation via the terms involving $\kappa_{j}$, $\gamma$, and $f$. The coupling terms involving $J_{2}$ and $J_{3}$ play a central role in generating nonreciprocity, with the dissipation modeled by the imaginary part of $J_{3}$ leading to asymmetric transmission. The full quantum dynamics, including these fluctuations, can be used to calculate observables such as the transmission coefficients, which reveal the nonreciprocal behavior of the system. These calculations are essential for understanding how energy dissipation and interference in the system give rise to unidirectional optical transmission. To solve Eq. (17), we perform the transformations $\delta a_{j}\rightarrow\delta\tilde{a}_{j}e^{-i\Delta^{\prime}_{j}t}$, $\delta d\rightarrow\delta\tilde{d}e^{-i\Delta_{\rm en}t}$, and $\delta b\rightarrow\delta\tilde{b}e^{-i\omega_{m}t}$. These transformations yield,